# Statistically Valid Post-Deployment Monitoring Should Be Standard for AI-Based Digital Health

### v1

@@ Abstract
This position paper argues that post-deployment monitoring in clinical AI is underdeveloped and proposes statistically valid and label-efficient testing frameworks as a principled foundation for ensuring reliability and safety in real-world deployment. A recent review found that only 9% of FDA-registered AI-based healthcare tools include a post-deployment surveillance plan muralidharanScopingReviewReporting2024 . Existing monitoring approaches are often manual, sporadic, and reactive, making them ill-suited for the dynamic environments in which clinical models operate. We contend that post-deployment monitoring should be grounded in label-efficient and statistically valid testing frameworks, offering a principled alternative to current practices. We use the term “statistically valid" to refer to methods that provide explicit guarantees on error rates (e.g., Type I/II error), enable formal inference under pre-defined assumptions, and support reproducibility—features that align with regulatory requirements. Specifically, we propose that the detection of changes in the data and model performance degradation should be framed as distinct statistical hypothesis testing problems. Grounding monitoring in statistical rigor ensures a reproducible and scientifically sound basis for maintaining the reliability of clinical AI systems. Importantly, it also opens new research directions for the technical community—spanning theory, methods, and tools for statistically principled detection, attribution, and mitigation of post-deployment model failures in real-world settings.
@@ 1 Introduction
AI models play a growing role in healthcare by providing advanced tools for disease diagnosis, medical imaging analysis, treatment planning, and patient monitoring de2018clinically ; mckinney2020international ; artificial2020state ; thorsen2020machine ; chenFrameworkIntegratingArtificial2023 . However, their promise is contingent on maintaining reliability and accuracy post-deployment—an area where significant challenges remain. AI-based digital health tools are known to experience performance degradation over time, which can have profound clinical implications, from missed diagnoses in radiology mckinney2020international to delayed interventions in critical care thorsen2020machine . These declines are particularly pronounced in diverse healthcare settings, where variations in demographics, deployment sites, and equipment can exacerbate diagnostic disparities obermeyerDissectingRacialBias2019 ; seyyed2021underdiagnosis ; zech2018variable ; kelly2019key , posing risks to patient safety.
@@
Despite these risks, a recent study revealed that only 9% of FDA-registered AI-based healthcare tools include a post-deployment surveillance plan muralidharanScopingReviewReporting2024 . Yet, the FDA’s guidelines for Software as a Medical Device (SaMD) fda2017samd emphasize the importance of ongoing model evaluation, including the use of prospective, statistically valid real-world performance monitoring to ensure continued safety, effectiveness, and performance. Similarly, the National Institute of Standards and Technology (NIST) emphasizes post-deployment monitoring as a cornerstone of its AI Risk Management Framework, essential for managing risk and maintaining trust throughout the AI lifecycle tabassiArtificialIntelligenceRisk2023 . In line with these expectations, it is critical to understand the mechanisms by which model performance can deteriorate after deployment in order to address the issue in a systematic and effective way.
@@
Performance degradation in deployed AI models can arise from various sources: shifts in patient demographics, evolution of clinical practices, changes in medical equipment or protocols, emergence of new disease patterns, and variations in data acquisition procedures seyyed2021underdiagnosis ; davis2017calibration ; zech2018variable ; wong2021quantification ; devries2023impact ; kelly2019key . At a high level, these data-related changes can be grouped into two categories: covariate shift shimodairaImprovingPredictiveInference2000 ; sugiyamaCovariateShiftAdaptation2007 and concept drift schlimmer1986incremental ; widmer1996learning ; Lu2019LearningUC . Covariate shift refers to changes in the input features, while the relationship between input features and labels remains unchanged. For example, shifts in patient demographics, alterations in data collection methods. On the other hand, concept drift refers to changes in the relationship between input features and labels, while the distribution of input features remains unchanged. This can occur due to shifts in clinical practice, new medical guidelines, changes in outcome prevalence, or the emergence of new confounding variables. Figure 1 (a) and (b) depict simple examples of covariate shift and concept drift, respectively, while Table 2 in the Appendix summarizes key causes and examples of both in the clinical AI.
@@
The described shifts in the data can make the model’s previously learned associations less accurate or outright invalid. Although model performance can be evaluated to determine whether these shifts change performance meaningfully, common evaluation methods after deployment are often based on average performance metrics performed manually and sporadically by clinicians davis2017calibration ; zech2018variable . By the time clinicians detect a decline in model performance, significant harm may have already occurred, and trust in the model may have been lost. Moreover, average performance metrics can mask degradation in specific patient subgroups kentOverallAverageTreatment2023 . Identifying and monitoring these subgroup-specific performance changes is crucial for ensuring effective care for all patients. However, effective and persistent post-deployment monitoring of this form is challenging due to the scarcity of ground truth labels berisha2024label ; manonmaniSemanticAnnotationHealthcare2021 ; xiaCorpusAnnotationChallenges2012 .
@@
While AI in healthcare includes both predictive and generative applications, this paper focuses exclusively on predictive models—such as those used for diagnosis, prognosis, and clinical risk scoring. We do not address generative models like large language models (LLMs). This focus is deliberate: even for predictive systems, post-deployment monitoring remains an unsolved challenge. Establishing rigorous methods in this domain is a necessary first step. This paper argues that post-deployment monitoring remains underemphasized in the machine learning community—particularly in high-stakes applications like clinical AI, where errors can have severe consequences. Current practices are ad hoc, sporadic, and reactive, lacking the systematic rigor needed to ensure safety and reliability. We contend that integrating statistically valid testing frameworks into post-deployment workflows offers a principled and label-efficient foundation, and should become a core component of the machine learning lifecycle for clinical applications.
@@ 2 Related Work
Covariate shift has been extensively studied in the machine learning community. Early theoretical work defined covariate shift and developed importance weighting techniques for adaptation shimodairaImprovingPredictiveInference2000 ; sugiyamaCovariateShiftAdaptation2007 . Subsequent research provided unified taxonomies of dataset shift types morenoTorresUnifyingViewDataset2012 and empirical studies evaluating drift detection methods in high-dimensional settings rabanser2019failing . With regard to concept drift, several detection methodologies have been proposed, reviews by luLearningConceptDrift2019 and gemaqueOverviewUnsupervisedDrift categorize drift detection techniques, including unsupervised settings where labels are scarce—a scenario common in clinical environments. In the context of our work, yuConceptDriftDetection2019 introduced a classifier-independent drift detector based on hierarchical hypothesis testing, one of the few existing approaches that aligns with the statistically principled framework we advocate.
@@
In healthcare applications, nestor2019feature demonstrated how shifts in patient demographics and clinical workflows could impact predictive models, while subbaswamy2021evaluating proposed practical evaluation strategies under distributional shifts, highlighting challenges particularly relevant for healthcare deployments. prathapanQuantifyingInputData2024 used CUSUM charts to track drift in medical AI, and zamzmiOutofDistributionDetectionRadiological2024 applied statistical process control for radiological data monitoring. Recent methods by davisDetectionCalibrationDrift2020 ; soinCheXstrayRealtimeMultiModal2022 proposed adaptive windowing strategies to capture gradual performance shifts.
@@
Finally, the operationalization of AI model monitoring has been advanced by the field of MLOps. While our focus is on statistically valid monitoring, there has been considerable development in system infrastructure. The work in khattakMLHOpsMachineLearning2023a proposed a comprehensive MLHOps framework, detailing deployment pipelines and monitoring components tailored for healthcare AI. Similarly, bhargavaChallengesSolutionsBest2024 identified architectural considerations and practical challenges for real-world post-deployment monitoring. This position paper complements these system-level frameworks by arguing for a statistically principled foundation for model monitoring, one that is label-efficient, interpretable, and aligned with regulatory expectations. We highlight how framing monitoring tasks as hypothesis testing problems enables systematic and actionable approaches that can be integrated into existing MLOps pipelines to enable safe deployment.
Figure 1: Framing Post-Deployment Monitoring as Hypothesis Testing. (a) Hypothesis tests for covariate shift, (b) concept drift. If a statistically significant change is observed, (c) a hypothesis test for model performance degradation is performed. If a model is affected by the change, (d) impacted subgroup identification is performed and used for target label collection and model recalibration. One of the open problems is addressing label scarcity for each of the described stages.
@@ 3 Framing Post-Deployment Monitoring as Hypothesis Testing
Given the challenges outlined above—manual monitoring, reliance on coarse average metrics, label scarcity, and FDA expectations for statistical validity—there is a need for rigorous and scalable tools in clinical AI. We propose reframing post-deployment monitoring as a collection of statistically grounded two-sample hypothesis tests. This framing enables formal decision-making with controlled error rates, moving beyond heuristic-based methods that lack statistical guarantees. It also provides a principled foundation for aligning model monitoring with regulatory standards and clinical safety needs.
To this end, we organize post-deployment monitoring into two core stages: (I) Data Shift Detection and (II) Model Performance Monitoring. Within the data shift stage, we introduce hypothesis tests for covariate shift (Section 4.1) and concept drift (Section 4.2). In the model performance stage, we develop tests for detecting degradation in overall model accuracy (Section 5.1) as well as distributional shifts in prediction correctness (Section 5.2). Each monitoring task is cast as a two-sample hypothesis test comparing pre- and post-deployment distributions, providing statistically grounded monitoring, as depicted in Figure 1.
While conceptually straightforward, this formulation exposes several unresolved challenges. Most notably, evaluating performance degradation requires access to post-deployment ground truth labels—which are often delayed, costly, or entirely unavailable. This motivates research into label-efficient monitoring strategies, including active learning, surrogate labeling, weak supervision, and model-based uncertainty estimation. Moreover, once degradation is detected, identifying the most affected subgroups remains an open problem critical to ensuring fairness and guiding retraining. Throughout the manuscript, we provide formal problem statements and examples of existing approaches and present these challenges as open problems (OP) for the community.
Let $\mathcal{D}_{t_{0}}=\{\mathbf{V}_{i}\}_{i=1}^{n_{0}}$ and $\mathcal{D}_{t_{1}}=\{\mathbf{U}_{i}\}_{i=1}^{n_{1}}$ denote i.i.d. samples collected at pre-deployment and post-deployment time points $t_{0}$ and $t_{1}$, drawn from distributions $p_{t_{0}}$ and $p_{t_{1}}$, respectively. Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a model mapping clinical inputs to predictions, and let $M_{t}=g(f,p_{t})$ denote its performance under distribution $p_{t}$. We frame monitoring tasks as two classes of two-sample tests:
@@
Reject $\text{Reject }H_{0}:$ $\text{if}\quad T(\mathcal{D}_{t_{0}},\mathcal{D}_{t_{1}})>c$ (3)
[…]
Table 1 summarizes candidate test statistics (both parametric and non-parametric) for these two-sample testing problems, along with trade-offs in assumptions and statistical power. We refer readers to Appendix B for the rationale and implementation details of these test statistics.
{tblr}
cell11 = c=2, cell21 = r=5, cell22 = r=3, cell52 = r=2, cell71 = r=6, cell74 = c=2, cell84 = c=2, cell92 = r=4, cell94 = c=2, cell104 = c=2, cell114 = c=2, cell124 = c=2, vlines, hline1-2,7 = -, hline3-4,6,10-12 = 3-5, hline5,8-9 = 2-5, hline7 = 2-, hline13 = -, & Tests When to Use / Notes Data Distribution Assumptions
@@
M    z test When population standard deviation is known normality, known variance
@@
V    F-Test Compare two variances; affected by skewness normality
@@
Non-Parametric M    Mann-Whitney U Test When distribution shape is unknown or non-normal
V    Levene’s Test When normality is uncertain
Distr. Shifts    Kolmogorov-Smirnov (KS) Test General-purpose test; best for moderate sample sizes
@@ 4 Data Shift Detection
One of the initial challenges in post-deployment monitoring is to detect data-only distributional changes between the pre- and post-deployment time points, denoted $t_{0}$ and $t_{1}$. We distinguish between two primary types of data shift: covariate shift, where the distribution of inputs changes while the input-output relationship remains fixed, and concept drift, where the conditional relationship between inputs and outputs changes, while the input distribution remains fixed morenoTorresUnifyingViewDataset2012 . Figure 1 (a) and (b) illustrate a simple example of covariate shift and data drift. In the following subsections, we formally define each type of shift and formulate hypothesis tests to detect them.
Let $\left(\mathbf{S},\mathbf{C}\right)$ denote a pair of random variables in the sample space $\mathcal{S}\times\mathcal{C}$, representing input and clinical feature variables, respectively. We separate $\mathbf{S}$ from $\mathbf{C}$ to reflect their distinct roles in the monitoring pipeline. The model operates solely on $\mathbf{S}$, which we assume encodes all relevant information for prediction, while $\mathbf{C}$ is treated as a latent variable that can be used in the identification of impacted subgroups. We use $Y\in\{0,1\}$ to denote the corresponding label random variable for $\left(\mathbf{S},\mathbf{C}\right)$, resulting in the tuple $\left(\mathbf{S},\mathbf{C},Y\right)$, which includes both feature and label variables. Assuming that $\left\{\left(\mathbf{S},\mathbf{C},Y\right)_{i}\right\}_{i=1}^{n}$ are i.i.d., we express the marginal distribution of $\left(\mathbf{S},\mathbf{C},Y\right)$ as $p\left(\mathbf{s},\mathbf{c},y\right)=p\left(y\mid\mathbf{s},\mathbf{c}\right)p\left(\mathbf{s},\mathbf{c}\right)$.

### v3

@@ Abstract
This position paper argues that post-deployment monitoring in clinical AI is underdeveloped and proposes statistically valid and label-efficient testing frameworks as a principled foundation for ensuring reliability and safety in real-world deployment. A recent review found that only 9% of FDA-registered AI-based healthcare tools include a post-deployment surveillance plan Muralidharan et al. (2024). Existing monitoring approaches are often manual, sporadic, and reactive, making them ill-suited for the dynamic environments in which clinical models operate. We contend that post-deployment monitoring should be grounded in label-efficient and statistically valid testing frameworks, offering a principled alternative to current practices. We use the term “statistically valid" to refer to methods that provide explicit guarantees on error rates (e.g., Type I/II error), enable formal inference under pre-defined assumptions, and support reproducibility—features that align with regulatory requirements. Specifically, we propose that the detection of changes in the data and model performance degradation should be framed as distinct statistical hypothesis testing problems. Grounding monitoring in statistical rigor ensures a reproducible and scientifically sound basis for maintaining the reliability of clinical AI systems. Importantly, it also opens new research directions for the technical community—spanning theory, methods, and tools for statistically principled detection, attribution, and mitigation of post-deployment model failures in real-world settings.
@@ 1 Introduction
AI models play a growing role in healthcare by providing advanced tools for disease diagnosis, medical imaging analysis, treatment planning, and patient monitoring De Fauw and others (2018); McKinney and others (2020); Benjamens et al. (2020); Thorsen-Meyer and others (2020); Chen et al. (2023). However, their promise is contingent on maintaining reliability and accuracy post-deployment—an area where significant challenges remain. AI-based digital health tools are known to experience performance degradation over time, which can have profound clinical implications, from missed diagnoses in radiology McKinney and others (2020) to delayed interventions in critical care Thorsen-Meyer and others (2020). These declines are particularly pronounced in diverse healthcare settings, where variations in demographics, deployment sites, and equipment can exacerbate diagnostic disparities Obermeyer et al. (2019); Seyyed-Kalantari and others (2021); Zech et al. (2018); Kelly et al. (2019), posing risks to patient safety.
@@
Despite these risks, a recent study revealed that only 9% of FDA-registered AI-based healthcare tools include a post-deployment surveillance plan Muralidharan et al. (2024). Yet, the FDA’s guidelines for Software as a Medical Device (SaMD) U.S. Food and Drug Administration (2017) emphasize the importance of ongoing model evaluation, including the use of prospective, statistically valid real-world performance monitoring to ensure continued safety, effectiveness, and performance. Similarly, the National Institute of Standards and Technology (NIST) emphasizes post-deployment monitoring as a cornerstone of its AI Risk Management Framework, essential for managing risk and maintaining trust throughout the AI lifecycle Tabassi (2023). In line with these expectations, it is critical to understand the mechanisms by which model performance can deteriorate after deployment in order to address the issue in a systematic and effective way.
@@
Performance degradation in deployed AI models can arise from various sources: shifts in patient demographics, evolution of clinical practices, changes in medical equipment or protocols, emergence of new disease patterns, and variations in data acquisition procedures Seyyed-Kalantari and others (2021); Davis et al. (2017); Zech et al. (2018); Wong et al. (2021); de Vries et al. (2023); Kelly et al. (2019). At a high level, these data-related changes can be grouped into three categories: covariate shift Shimodaira (2000); Sugiyama et al. (2007), label shift Garg et al. (2020), and concept drift Schlimmer and Granger Jr (1986); Widmer and Kubat (1996); Lu et al. (2019). Covariate shift refers to changes in the input features while the relationship between the input features and the labels remains unchanged. For example, shifts in patient demographics and alterations in data collection methods. Label shift occurs due to the changes in the output distribution, while output-conditional feature distributions remain the same. For example, seasonal fluctuations in flu prevalence. On the other hand, concept drift refers to changes in the relationship between input features and labels, while the distribution of input features remains unchanged. This can occur due to shifts in clinical practice, new medical guidelines, changes in outcome prevalence, or the emergence of new confounding variables. Figure 1 (a), (b), and (c) depict simple examples of covariate shift, label shift, and concept drift, respectively, while Table 2 in the Appendix A summarizes key causes and examples in the clinical AI.
@@
The described shifts in the data can make the model’s previously learned associations less accurate or outright invalid. Although model performance can be evaluated to determine whether these shifts change performance meaningfully, common evaluation methods after deployment are often based on average performance metrics performed manually and sporadically by clinicians Davis et al. (2017); Zech et al. (2018). By the time clinicians detect a decline in model performance, significant harm may have already occurred, and trust in the model may have been lost. Moreover, average performance metrics can mask degradation in specific patient subgroups Kent (2023). Identifying and monitoring these subgroup-specific performance changes is crucial for ensuring effective care for all patients. However, effective and persistent post-deployment monitoring of this form is challenging due to the scarcity of ground truth labels Li et al. (2024b); Manonmani and Balakrishanan (2021); Xia and Yetisgen-Yildiz (2012).
@@
While AI in healthcare includes both predictive and generative applications, this paper focuses exclusively on predictive models—such as those used for diagnosis, prognosis, and clinical risk scoring. We do not address generative models like large language models (LLMs). This focus is deliberate: even for predictive systems, post-deployment monitoring remains an unsolved challenge. Establishing rigorous methods in this domain is a necessary first step. This paper argues that post-deployment monitoring remains underemphasized in the machine learning community—particularly in high-stakes applications like clinical AI, where errors can have severe consequences. Current practices are ad hoc, sporadic, and reactive, lacking the systematic rigor needed to ensure safety and reliability. We contend that integrating statistically valid testing frameworks into post-deployment workflows offers a principled and label-efficient foundation and should become a core component of the machine learning lifecycle for clinical applications.
@@ 2 Related Work
Covariate shift has been extensively studied in the machine learning community. Early theoretical work defined covariate shift and developed importance weighting techniques for adaptation Shimodaira (2000); Sugiyama et al. (2007). Subsequent research provided unified taxonomies of dataset shift types Moreno-Torres et al. (2012) and empirical studies evaluating drift detection methods in high-dimensional settings Rabanser et al. (2019). With respect to concept drift, Lu et al. (2019) offers a comprehensive taxonomy of drift types and adaptation strategies in data-stream learning, focusing primarily on supervised and semi-supervised settings. Gemaque et al. (2020) focuses on unsupervised scenarios where labels are scarce and categorizes detectors based on statistical, clustering, and reconstruction principles. In the context of our work, Yu et al. (2019) introduces a classifier-independent drift detector based on hierarchical hypothesis testing, one of the few existing approaches that aligns with the statistically principled framework we advocate. Finlayson et al. (2021) highlights the risks of dataset shift in deployed clinical ML systems and advocates for clinician-in-the-loop monitoring to detect and respond to data shifts, emphasizing governance and oversight.
A range of statistical tests can be applied to compare pre- and post-deployment distributions, enabling formal detection of a shift through two-sample hypothesis testing. Table 1 summarizes common parametric and nonparametric tests, which can be applied to both data-shift detection Rabanser et al. (2019); Gretton et al. (2012) and model-performance monitoring Davis et al. (2020a).
@@
At deployment scale, many concurrent monitors (features, subgroups, metrics) benefit from online multiple-testing with anytime-valid FDR control (SAFFRON) and e-value procedures (e-BH) for adaptive false-alarm control (Ramdas et al., 2018; Wang and Ramdas, 2022). In healthcare applications, Nestor et al. (2019) demonstrated how shifts in patient demographics and clinical workflows can degrade predictive-model performance, while Subbaswamy et al. (2021) proposed practical evaluation strategies under distributional shift, highlighting challenges particularly relevant to clinical deployments. Prathapan et al. (2024) used CUSUM control charts to track input drift in medical AI, and Zamzmi et al. (2024) applied statistical process control methods for radiological data monitoring. Davis et al. (2020b) focused on detecting calibration drift in predictive models, whereas Soin et al. (2022) introduced adaptive windowing for real-time multimodal performance monitoring.
@@
Finally, the operationalization of AI model monitoring has been advanced by the field of MLOps. While our focus is on statistically valid monitoring, there has been considerable development in system infrastructure. The work in Khattak et al. (2023) proposed a comprehensive MLHOps framework, detailing deployment pipelines and monitoring components tailored for healthcare AI. Similarly, Bhargava and Singhal (2024) identified architectural considerations and practical challenges for real-world post-deployment monitoring. This position paper complements these system-level frameworks by arguing for a statistically principled foundation for model monitoring, one that is label-efficient, interpretable, and aligned with regulatory expectations. We highlight how framing monitoring tasks as hypothesis testing problems enables systematic and actionable approaches that can be integrated into existing MLOps pipelines to enable safe deployment.
Figure 1: Framing Post-Deployment Monitoring as Hypothesis Testing. Binary classification example. A hypothesis test for (a) covariate shift—input features distribution changes, but the relationship between labels and input features remains the same, (b) label shift—output distribution changes, while output-conditional feature distributions remain the same, (c) concept drift—relationship between labels and input features changes, but the distribution of the input features remains the same. If a statistically significant change is observed, a hypothesis test for (d) model performance degradation is performed. If a model is affected by the change, (e) impacted subgroup identification is performed and used for target label collection and model recalibration. One of the open problems is addressing label scarcity for each of the described stages.
@@ 3 Framing Post-Deployment Monitoring as Hypothesis Testing
Given the challenges outlined in the introduction—manual monitoring, reliance on coarse average metrics, label scarcity, and FDA expectations for statistical validity—there is a need for rigorous and scalable monitoring tools in clinical AI. We propose reframing post-deployment monitoring as a collection of statistically grounded two-sample hypothesis tests. This framing enables formal decision-making with controlled error rates, moving beyond heuristic-based methods that lack statistical guarantees. It also provides a principled foundation for aligning model monitoring with regulatory standards and clinical safety needs.
To this end, we organize post-deployment monitoring into two core stages: (I) Data Shift Detection and (II) Model Performance Monitoring. Within the data shift stage, we introduce hypothesis tests for covariate shift (Section 4.1) and concept drift (Section 4.2). In the model performance stage, we develop tests for detecting degradation in overall model accuracy (Section 5.1) as well as distributional shifts in prediction correctness - e.g. do errors concentrate in specific subgroups? (Section 5.2). Each monitoring task is cast as a two-sample hypothesis test comparing pre- and post-deployment distributions, providing statistically grounded monitoring, as depicted in Figure 1. The post-deployment monitoring is modality-agnostic, as long as the data can be represented as input features and clinical/demographic variables.
While conceptually straightforward, this formulation exposes several unresolved challenges. Most notably, evaluating performance degradation requires access to post-deployment ground truth labels, which are often delayed, costly, or entirely unavailable. In high-dimensional settings, required sample sizes rise and test power can drop; thus, label-efficient tactics (e.g., sequential looks and active labeling) and explicit reporting of decision margins/thresholds are essential. This motivates research into label-efficient monitoring strategies, including active learning, surrogate labeling, weak supervision, and model-based uncertainty estimation. Moreover, once degradation is detected, identifying the most affected subgroups remains an open problem critical to ensuring fairness and guiding retraining. Throughout the manuscript, we provide formal problem statements and examples of existing approaches and present these challenges as open problems (OP) for the community.
Let $\mathcal{D}_{t_{0}}$ and $\mathcal{D}_{t_{1}}$ denote i.i.d. samples 111i.i.d. assumption is local within each period (samples at $t_{0}$ or $t_{1}$) but not across time. collected at pre-deployment and post-deployment time points $t_{0}$ and $t_{1}$, drawn from distributions $p_{t_{0}}$ and $p_{t_{1}}$, respectively. Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a model mapping clinical inputs to predictions, and let $M_{t}=g(f,p_{t})$ denote its performance under distribution $p_{t}$. We frame monitoring tasks as two classes of two-sample tests:
@@
$\text{Reject }H_{0}:$ $\text{if}\quad T(\mathcal{D}_{t_{0}},\mathcal{D}_{t_{1}})>c$ (3)
$\text{Accept }H_{0}:$ $\text{otherwise}.$
Table 1 summarizes candidate two-sample test statistics—parametric and non-parametric—and the associated assumptions and power trade-offs. Appendix D outlines our test-selection rationale, and Appendices F and G describe the parametric and non-parametric tests, respectively.
Tests When to Use / Notes Data Distribution Assumptions
@@
M
   z test When population standard deviation is known normality, known variance
@@
V
   F-Test Compare two variances; affected by skewness normality
@@
Non-Parametric
M
   Mann-Whitney U Test When distribution shape is unknown or non-normal
V
   Levene’s Test When normality is uncertain
Distr. Shifts
   Kolmogorov-Smirnov (KS) Test General-purpose test; best for moderate sample sizes
@@ 4 Data Shift Detection
One of the initial challenges in post-deployment monitoring is to detect data-only distributional changes between the pre- and post-deployment time points, denoted $t_{0}$ and $t_{1}$. We distinguish between three types of data shift: covariate shift, where the distribution of inputs changes while the input-output relationship remains fixed, label shift, where the marginal distribution of outputs changes while the distribution of inputs given outputs remains fixed; and concept drift, where the conditional relationship between inputs and outputs changes, while the input distribution remains fixed Moreno-Torres et al. (2012). Figure 1 (a), (b), and (c) illustrate a simple example of covariate shift, label shift, and concept drift. In the following subsections, we define covariate shift and concept drift and develop hypothesis tests to detect them. We defer label shift to Appendix B, since its testing procedure is analogous to the covariate shift test.
Let $\left(\mathbf{S},\mathbf{C}\right)$ denote a pair of random variables in the sample space $\mathcal{S}\times\mathcal{C}$, representing input and clinical feature variables, respectively. We separate $\mathbf{S}$ from $\mathbf{C}$ to reflect their distinct roles in the monitoring pipeline. The model operates solely on $\mathbf{S}$, which we assume encodes all relevant information for prediction, while $\mathbf{C}$ is treated as a latent variable that can be used in the identification of impacted subgroups. We use $Y\in\{0,1\}$ to denote the corresponding label random variable for $\left(\mathbf{S},\mathbf{C}\right)$, resulting in the tuple $\left(\mathbf{S},\mathbf{C},Y\right)$, which includes both feature and label variables. Assuming that $\left\{\left(\mathbf{S},\mathbf{C},Y\right)_{i}\right\}_{i=1}^{n}$ are i.i.d., we express the marginal distribution of $\left(\mathbf{S},\mathbf{C},Y\right)$ as $p\left(\mathbf{s},\mathbf{c},y\right)=p\left(y\mid\mathbf{s},\mathbf{c}\right)p\left(\mathbf{s},\mathbf{c}\right)$, where $\mathbf{s}$, $\mathbf{c}$ and $y$ denote the realizations of the random variables.